# Block-to-Scene Pre-training for Point Cloud Hybrid-Domain Masked Autoencoders

### v1

@@ Abstract
Point clouds, as a primary representation of 3D data, can be categorized into scene domain point clouds and object domain point clouds based on the modeled content. Masked autoencoders (MAE) have become the mainstream paradigm in point clouds self-supervised learning. However, existing MAE-based methods are domain-specific, limiting the model’s generalization. In this paper, we propose to pre-train a general Point cloud Hybrid-Domain Masked AutoEncoder (PointHDMAE) via a block-to-scene pre-training strategy. We first propose a hybrid-domain masked autoencoder consisting of an encoder and decoder belonging to the scene domain and object domain, respectively. The object domain encoder specializes in handling object point clouds and multiple shared object encoders assist the scene domain encoder in analyzing the scene point clouds. Furthermore, we propose a block-to-scene strategy to pre-train our hybrid-domain model. Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space. Then, we employ an object-level mask and reconstruction pipeline to recover the masked points of each block, enabling the object encoder to learn a universal object representation. Finally, we introduce a scene-level block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, facilitating the learning of scene representations. Extensive experiments across different datasets and tasks demonstrate the generalization and superiority of our hybrid-domain model. The code will be released.
@@ 1 Introduction
With the rapid development of 3D scanning technology, 3D point clouds have become the mainstream representation for 3D objects due to their ease of acquisition, explicit representation, and efficient storage. Point clouds can be categorized into scene domain point clouds (Dai et al., 2017; Song et al., 2015; Armeni et al., 2016; Zheng et al., 2020; Sun et al., 2020) and object domain point clouds (Wu et al., 2015; Chang et al., 2015; Uy et al., 2019; Deitke et al., 2024; Yu et al., 2023) based on the modeling object. As shown in Figure 1 (a), object domain point clouds describe specific objects or entities, such as an airplane, with relatively fewer points. Scene domain point clouds represent the entire environment or scene, such as indoor scenes, including multiple objects, structures, and background elements, with a larger number of points. Due to the significant disparity in point count and the elements being described, a substantial domain gap exists in these two types of point clouds.
@@
Figure 1: The handling approach for point clouds from different domains in domain-specific models (a) compared to our hybrid-domain model (b), and the generalization experiments of the domain-specific Point-MAE model (c).
@@
Recently, point cloud masked autoencoders (Yu et al., 2022; Pang et al., 2022; Zhang et al., 2022; Dong et al., 2023; Zha et al., 2024a), pre-trained on massive point cloud data, have become the mainstream paradigm in point cloud self-supervised learning and have been widely applied to various point cloud tasks. It is inspired by masked image modeling (Bao et al., 2021; He et al., 2022; Xie et al., 2022), using the unmasked portions to predict the geometric coordinates or semantic features of the masked parts, thereby enabling the model to learn universal 3D representations. Despite the significant success, most of these methods are domain-specific due to the notable domain gap between object-domain and scene-domain point clouds. As shown in Figure 1 (a), these methods (Yu et al., 2022; He et al., 2022; Zhang et al., 2022; Zha et al., 2023b; Zhou et al., 2024) use scene-level models for scene tasks and object-level models for object tasks, thereby limiting their generalizability. For example, Point-MAE (Pang et al., 2022), which is pre-trained on ShapeNet (Chang et al., 2015), primarily performs object point cloud tasks. For scene point clouds, it requires re-pretraining on scene-level datasets like ScanNet (Dai et al., 2017) to adapt to the scene domain. As shown in Figure 1 (c), directly transferring an object domain pre-trained model (e.g.Point-MAE (Pang et al., 2022)) to scene tasks results in a significant performance drop. Similarly, models pre-trained on the scene domain also exhibit performance declines when transferred to object task.
@@
Pre-training a general point cloud model is our persistent pursuit; however, it is highly challenging for two main reasons. Firstly, the input data is inconsistent. Scene-level point clouds, such as ScanNet (Dai et al., 2017), typically consist of 50k points, while object-level point clouds like ModelNet (Wu et al., 2015) typically consist of 1k points. The disparity in point count between the two types is significant, making it difficult to process both types of data simultaneously using a single model. Secondly, there is inconsistency in task emphasis. Scene point clouds typically involve object detection or segmentation, which often prioritizes understanding fine-grained local point clouds. On the contrary, object point clouds generally involve classification tasks, which tend to prioritize understanding global geometry.
@@
To address the aforementioned challenges, we propose a block-to-scene pretraining strategy to pre-train a Point cloud Hybrid-Domain Masked Auto-Encoder (PointHDMAE). We address the challenge of inconsistent input data by using domain-specific encoders to process data from their respective domains. Additionally, we finetune the pre-trained model to address the inconsistency of task emphasis. Specifically, as shown in Figure 2, we first design a point cloud hybrid-domain architecture that consists of an encoder and decoder belonging to the scene domain and object domain, respectively. In the fine-tuning phase, as shown in Figure 1 (b), for object domain data, our model selectively activates the object-domain encoder for analysis. However, in the case of scene point clouds, we activate multiple shared object encoders to assist the scene encoder in analyzing scene domain data collaboratively.
@@
Furthermore, we propose a block-to-scene pre-training strategy that couples masked reconstruction and position regression tasks of random object blocks within a scene for self-supervised learning, enabling us to train encoders for different domains simultaneously. Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space. Then, within the object domain, we use a mask and reconstruction pipeline to recover the masked points of each block, enabling it to learn universal object representations. Finally, we introduce a scene-level block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, enabling the scene encoder to learn scene representations with the assistance of the object encoders. By block-to-scene pretraining, our model can simultaneously learn powerful object-level and scene-level representations and exhibit superior transferability. Our model can be fine-tuned directly on downstream tasks such as object point cloud classification, segmentation, completion, and scene point cloud detection without the need for any additional domain adaptation training.
Our main contributions can be summarized as follows: (1) We propose a point cloud hybrid-domain masked autoencoders to address the generalization limitations of existing domain-specific MAE-based models. (2) We propose a block-to-scene pretraining strategy, a joint pre-training strategy that reconstructs and regresses random object blocks within a scene. (3) Extensive experiments across different datasets and tasks demonstrate the generalization and superiority of our model.
@@ 2 Related Work
2.1 Self-supervised Learning for Point Cloud.
Self-supervised learning, which enables the learning of general representations from large amounts of unlabeled data, has been widely applied in fields such as language (Semnani et al., 2019; Brown et al., 2020; Achiam et al., 2023) and image (Bao et al., 2021; Chen et al., 2020b; a; He et al., 2022). Inspired by the success of visual pretraining, numerous point cloud pretraining methods have also been proposed. Based on the pretext tasks, they can be categorized into contrastive learning paradigms (Oord et al., 2018; Tian et al., 2020) and masked reconstruction paradigms (Bao et al., 2021; He et al., 2022). PointContrast (Xie et al., 2020b), CrossPoint (Afham et al., 2022), and DepthContrast (Zhang et al., 2021) construct positive and negative sample pairs using various methods and employ contrastive learning techniques to learn 3D representations.
Point-BERT (Yu et al., 2022) was the first to propose learning universal 3D representations using the paradigm of masked reconstruction. Subsequently, numerous explorations (Pang et al., 2022; Zhang et al., 2022; Zha et al., 2024a; b) have improved masked reconstruction from various perspectives. Point-MAE (Pang et al., 2022) and PointFEMAE (Zha et al., 2024a) introduced the masked autoencoder for reconstruction, and PointGPT (Chen et al., 2024) proposed pretraining using an autoregressive approach. To address the limited amount of point cloud during pretraining, many approaches integrate multimodal knowledge to aid in learning point cloud features. ACT (Dong et al., 2023) leverages a pre-trained image model to assist in point cloud reconstruction, while I2P-MAE (Zhang et al., 2023) employs an image-guided masking strategy. PiMAE (Chen et al., 2023) proposes to address the challenges of multi-modal interaction between point cloud and RGB image data through mask alignment, a two-branch MAE pipeline, and a cross-modal reconstruction module. In this paper, we propose a block-to-scene pretraining strategy that combines masked reconstruction and position regression in a joint self-supervised learning method to enhance the model’s generalizability.
@@ 3 Methodology
In this section, we provide a detailed explanation of how to use our block-to-scene pretraining strategy to train our point cloud hybrid-domain masked autoencoder.
3.1 Point Cloud Hybrid-Domain Masked Autoencoder (PointHDMAE)
The overall architecture of our PointHDMAE is shown in Figure 2, it is composed of four main components: a scene encoder, a scene decoder, a shared object encoder, and a shared object decoder. It is primarily used for two main task pipelines: object point cloud processing and scene point cloud processing.
Our PointHDMAE is a hybrid model. Due to significant differences across domains of point clouds and tasks, we selectively activate different sub-modules for various downstream data and tasks. For instance, in tasks such as object point cloud classification and object part segmentation, we selectively activate our object encoder according to specific tasks. For scene point cloud detection tasks, we activate all encoders. This collaborative approach is primarily adopted because utilizing the object encoder for analyzing local point blocks within the scene contributes to the scene encoder’s comprehension of scene intricacies.
@@ 3.2 Block-to-Scene Pretraining
Figure 2: The architecture of our point cloud hybrid-domain masked autoencoder and the pipeline for block-to-scene pre-training. The left side illustrates the scene-level block position regression, while the right side shows the object-level block masked reconstruction.
@@ 3.2.1 Random Point Block Generation
Ground-truth block position generation. We generate the ground truth position of each random point block in the scene, which will be used to constrain the predicted position regressed by the final scene decoder. Inspired by the detection (Carion et al., 2020; Misra et al., 2021; Dai et al., 2021) task, we use the 3D bounding box of each random point block as its ground truth position. By computing the mean of all points in each dimension of the entire point block, the coordinates of the center point are obtained. The half-lengths of the bounding box in each dimension are calculated by subtracting the minimum value from the maximum value in each dimension and dividing by 2. Subsequently, the center point coordinates and half-lengths in each dimension (x, y, and z) are concatenated to form the bounding box. Finally, standard procedures (Misra et al., 2021) are applied to compute bounding box parameters $\bm{B_{b}}$ such as size and corners for each bounding box.
@@ 3.2.2 Object-level Block Masked Reconstruction
Object point block generation. We treat each randomly selected point block in $\bm{B}$ as an object point block and transform its coordinates from the scene space to the object space for processing by the object autoencoder. Specifically, we apply a simple set of transformation functions to each point block. First, we subtract the coordinates of the center point from the $N_{o}$ local points to obtain the relative coordinates of each point. Then, we normalize these coordinates to the range [-1, 1]. Finally, after applying a random rotation transformation to each point block, we obtain all point blocks $\bm{B_{o}=\{B_{o}^{1},...,B_{o}^{K_{o}}\}}$ as input to the object encoder. Through these transformation functions, we convert the coordinates of each point block from the scene space to the object space, decoupling the point block object coordinates from the original scene coordinates. This enables the object encoder to learn the universal shape features of the point block objects.
Object point block masked and reconstruction. We use a shared object autoencoder to perform mask-based reconstruction self-supervised learning on all generated object point blocks $\bm{B_{o}}$, enabling our object encoder to learn a general representation of objects. We illustrate the entire mask and reconstruction process for the object point blocks using Point-MAE (Pang et al., 2022) as an example. For the $i$-th object point block $\bm{B_{o}^{i}\in\mathbb{R}^{N_{o}\times 3}}$, we use farthest point sampling and the K-nearest neighbors algorithm to divide it into $M_{o}$ point patches. Then, after randomly masking most of the patches, we generate initial tokens and positional encodings for each unmasked patch using MLP-based token encoding and positional encoding. By adding them, we obtain the token $\bm{E}_{o}^{0}\in\mathbb{R}^{rM_{o}\times C_{o}}$ for each unmasked patch, where $r$ represents the unmasked ratio, and $C_{o}$ denotes the object feature dimension. Finally, we use a shared object encoder to extract object features $\bm{E}_{o}^{n_{o}}\in\mathbb{R}^{rM_{o}\times C_{o}}$, where $n_{o}$ is the number of layers in the scene encoder.
@@ 3.2.3 Scene-level Block Position Regression
Scene encoding. Given an input point cloud $\bm{P_{s}}\in\mathbb{R}^{N_{s}\times 3}$ with $N_{s}$ points, we first use farthest point sampling and the K-nearest neighbors algorithm to partition it into blocks. Then, using an MLP-based token encoding layer and a positional encoding layer, we generate the semantic token and positional encoding for each patch. By adding them, we obtain the token $\bm{E}_{s}^{0}\in\mathbb{R}^{M_{s}\times C_{s}}$ for each patch, where $M_{s}$ represents the number of scene patches, and $C_{s}$ denotes the scene feature dimension. Finally, we use a scene encoder based on the standard Transformer (Vaswani et al., 2017) architecture to extract scene features $\bm{E}_{s}^{n_{s}}\in\mathbb{R}^{M_{s}\times C_{s}}$, where $n_{s}$ is the number of Transformer layers in the scene encoder.
Scene decoding and position regression. We apply max pooling to the features of all blocks output by the object encoder to obtain the global feature for each block. After passing through the projection layer, these point block features are transformed into features $\bm{B}_{g}\in\mathbb{R}^{K_{o}\times C_{s}}$ for scene decoding input. However, we prevent the gradients of $\bm{B}_{g}$ from propagating backward into the mask reconstruction pipeline during the backpropagation process, thereby mitigating the multi-task interference caused by the scene regression task on the object encoder. We will provide a detailed explanation of this issue in Section 4.3.1. We then add the transformed point block features to randomly initialized queries to obtain enhanced queries $\bm{Q^{0}}\in\mathbb{R}^{q\times C_{s}}$, where $q$ is the number of queries. Since the number of point blocks and queries often differ, we replicate $\bm{B}_{g}$ to match all queries.
We use a Transformer decoder based on self-attention and cross-attention as our scene decoder. The input queries $\bm{Q^{0}}$, with the assistance of the encoded features $\bm{E}_{s}^{n_{s}}$, pass through our decoder to obtain the decoded query features $\bm{Q^{m_{s}}}$, where $m_{s}$ is the number of scene decoder. Finally, we use different MLP-based reconstruction heads to predict the 3D bounding box $\bm{B_{b}^{p}}$ of each random point block.
@@
We use a combination of mask reconstruction loss and point block regression loss to jointly constrain our pre-training process. For the reconstruction loss calculation, we follow previous work (Pang et al., 2022; Zhang et al., 2022) and use Chamfer Distance (Fan et al., 2017) (CD) as the loss function. For the regression loss calculation, we reference detection (Misra et al., 2021) and use Generalized Intersection over Union (Rezatofighi et al., 2019) (GIoU) as the loss function. Therefore, our loss function is defined as follows:

### v3

# Point Cloud Mixture-of-Domain-Experts Model for 3D Self-supervised Learning

@@ Abstract
Point clouds, as a primary representation of 3D data, can be categorized into scene domain point clouds and object domain point clouds. Point cloud self-supervised learning (SSL) has become a mainstream paradigm for learning 3D representations. However, existing point cloud SSL primarily focuses on learning domain-specific 3D representations within a single domain, neglecting the complementary nature of cross-domain knowledge, which limits the learning of 3D representations. In this paper, we propose to learn a comprehensive Point cloud Mixture-of-Domain-Experts model (Point-MoDE) via a block-to-scene pre-training strategy. Specifically, We first propose a mixture-of-domain-expert model consisting of scene domain experts and multiple shared object domain experts. Furthermore, we propose a block-to-scene pretraining strategy, which leverages the features of point blocks in the object domain to regress their initial positions in the scene domain through object-level block mask reconstruction and scene-level block position regression. By integrating the complementary knowledge between object and scene, this strategy simultaneously facilitates the learning of both object-domain and scene-domain representations, leading to a more comprehensive 3D representation. Extensive experiments in downstream tasks demonstrate the superiority of our model.
@@ 1 Introduction
Recently, with the rapid development of 3D scanning technology, 3D point clouds have become the mainstream representation for 3D objects due to their ease of acquisition, explicit representation, and efficient storage. Point clouds can be categorized into scene domain point clouds Dai et al. (2017); Song et al. (2015); Armeni et al. (2016) and object domain point clouds Wu et al. (2015); Chang et al. (2015); Uy et al. (2019) based on the modeling object. As shown in Figure 2 (a), object domain point clouds describe specific objects or entities, such as an airplane, with relatively fewer points. Scene domain point clouds represent the entire environment or scene, such as indoor scenes, including multiple objects, structures, and background elements, with a larger number of points. Due to the significant disparity in point count and the elements being described, a substantial domain gap exists in these two types of point clouds.
@@
Figure 1: Comparison of (a) the traditional domain-specific model and (b) our mixture-of-domain-experts model for processing point clouds from different domains.
@@
Point cloud self-supervised learning (SSL) Yu et al. (2022); Pang et al. (2022); Zhang et al. (2022); Dong et al. (2023), pre-trained on massive point cloud data, have become the mainstream paradigm for learning 3D representations and have been widely transferred to various point cloud tasks Zhou et al. (2024); Zha et al. (2025). Despite the significant success, most of these methods are domain-specific due to the notable domain gap between object-domain and scene-domain point clouds. As shown in Figure 2 (a), these methods Yu et al. (2022); He et al. (2022); Zhang et al. (2022) use scene-level data to learn scene model and object-level data for object model. This domain-independent learning approach not only fails to effectively integrate knowledge across different domains, but also struggles to perform well on data from multiple domains simultaneously. For example, Point-MAE Pang et al. (2022), which is pre-trained on ShapeNet Chang et al. (2015), primarily performs object point cloud tasks. For scene point clouds, it requires re-pretraining on scene-level datasets like ScanNet Dai et al. (2017) to adapt to the scene domain. As shown in Figure 2 (c), directly transferring an object domain Point-MAE to scene tasks results in a significant performance drop. Similarly, models pre-trained on the scene domain also exhibit performance declines when transferred to object task.
Figure 2: Transferability of the domain-specific Point-MAE model. Transferring a Point-MAE pre-trained on the scene to an object task results in a clear performance drop, and the same occurs for models in the object domain..
@@
Learning a more comprehensive 3D representation by simultaneously integrating knowledge from both the object domain and the scene domain is a reasonable solution. However, it is highly challenging for two main reasons. Firstly, the input data is inconsistent. Scene-level point clouds, such as ScanNet Dai et al. (2017), typically consist of 50k points, while object-level point clouds like ModelNet Wu et al. (2015) typically consist of 1k points. The disparity in point count between the two types is significant, making it difficult to process both types of data simultaneously using a single model. Secondly, there is inconsistency in task emphasis. Scene point clouds typically involve object detection or segmentation, which often prioritizes understanding fine-grained local point clouds. On the contrary, object point clouds generally involve classification tasks, which tend to prioritize understanding global geometry.
@@
To address the aforementioned challenges, we propose a block-to-scene pretraining strategy to pre-train a Point cloud Mixture-of-Domain-Experts model (Point-MoDE). We address the challenge of inconsistent input data by using domain-specific experts to process data from their respective domains. Additionally, we finetune the pre-trained model to address the inconsistency of task emphasis. Specifically, as shown in Figure 2 (b), we first design a point cloud mixture-of-domain-experts model that consists of a scene domain expert and multiple shared object domain experts. In the fine-tuning phase, for object domain data, our model selectively activates the object-domain expert for analysis. However, in the case of scene point clouds, we activate multiple shared object expert encoders to assist the scene expert in analyzing scene domain data collaboratively.
@@
Furthermore, we propose a block-to-scene pre-training strategy that couples masked reconstruction and position regression tasks of random object blocks within a scene for self-supervised training, enabling us to train experts for different domains simultaneously. Specifically, we first randomly select point blocks within a scene and apply a set of transformations to convert each point block coordinates from the scene space to the object space. Then, within the object expert, we use a mask and reconstruction pipeline to recover the masked points of each block, enabling it to learn universal object representations. Finally, we introduce a scene-expert-based block position regression pipeline, which utilizes the blocks’ features in the object space to regress these blocks’ initial positions within the scene space, enabling the scene expert to learn scene representations with the assistance of the object experts. By block-to-scene pretraining, our Point-MoDE can simultaneously learn powerful object-level and scene-level representations and exhibit superior transferability. Our model can be fine-tuned directly on downstream tasks such as object point cloud classification, segmentation, completion, and scene point cloud detection without the need for any additional domain adaptation training.
The main contributions can be summarized as follows:
  • We propose a mixture-of-domain-experts model, integrating both scene domain and object domain experts to effectively process point clouds from different domains.
  • We introduce a novel block-to-scene pretraining strategy that combines masked reconstruction and position regression tasks, enabling the simultaneous learning of object-level and scene-level representations.
  • Extensive experiments across different datasets and tasks demonstrate the superiority and transferability of our model.
Figure 3: The architecture of our point cloud mixture-of-domain-experts model and the pipeline for block-to-scene pre-training. The left side illustrates the scene-level block position regression, while the right side shows the object-level block masked reconstruction.
@@ 2 Related Work
2.1 Self-supervised Learning for Point Cloud
Self-supervised learning, which enables the learning of general representations from large amounts of unlabeled data, has been widely applied in fields such as language Semnani et al. (2019); Brown et al. (2020); Achiam et al. (2023) and image Bao et al. (2021); Chen et al. (2020b, a); He et al. (2022). Inspired by the success of visual pretraining, numerous point cloud pretraining methods have also been proposed. Based on the pretext tasks, they can be categorized into contrastive learning paradigms Oord et al. (2018); Tian et al. (2020) and masked reconstruction paradigms Bao et al. (2021); He et al. (2022). PointContrast Xie et al. (2020), CrossPoint Afham et al. (2022), and DepthContrast Zhang et al. (2021) construct positive and negative sample pairs using various methods and employ contrastive learning techniques to learn 3D representations. Point-BERT Yu et al. (2022) was the first to propose learning universal 3D representations using the paradigm of masked reconstruction. Subsequently, numerous explorations have improved masked reconstruction from various perspectives. Point-MAE Pang et al. (2022) and Point-M2AE Zhang et al. (2022) introduced the masked autoencoder for reconstruction, and PointGPT Chen et al. (2024) proposed pretraining using an autoregressive approach. To address the limited amount of point cloud during pretraining, many approaches integrate multimodal knowledge to aid in learning point cloud features. ACT Dong et al. (2023) leverages a pre-trained image model to assist in point cloud reconstruction, while I2P-MAE Zhang et al. (2023) employs an image-guided masking strategy. Despite the significant success, the above methods are domain-specific. In this paper, we propose training a mixture-of-domain-experts model to integrate knowledge from different domains for learning comprehensive 3D representations.
@@ 3 Methodology
In this section, we provide a detailed explanation of how to use our block-to-scene pretraining strategy to train our point cloud mixture-of-domain-experts model.
3.1 Point Cloud Mixture-of-Domain-Experts
The overall architecture of our Point-MoDE is shown in Figure 3, it is composed of four main components: a scene expert encoder, a scene expert decoder, a shared object expert encoder, and a shared object expert decoder. It is primarily used for two main task pipelines: object point cloud processing and scene point cloud processing.
Our Point-MoDE is a hybrid model. Due to significant differences across domains of point clouds and tasks, we selectively activate different sub-modules for various downstream data and tasks. For instance, in tasks such as object point cloud classification and object part segmentation, we selectively activate our object expert encoder according to specific tasks. For scene point cloud detection tasks, we activate all encoders. This collaborative approach is primarily adopted because utilizing the object expert encoder for analyzing local point blocks within the scene contributes to the scene encoder’s comprehension of scene intricacies.
In the scene expert model, we adopt a standard Transformer as our scene expert, comprising 3 layers of Transformer-based encoders and 8 layers of Transformer-based decoders with a PointNet-based Qi et al. (2017a) token embedding layer. In our object expert model, we directly leverage existing pre-trained object models based on mask-based point modeling, such as Point-BERT Yu et al. (2022), Point-MAE Pang et al. (2022), PointGPT Chen et al. (2024), etc. Using these models pre-trained on the object domain allows us to incorporate object-level priors into our pipeline while saving training resources and enabling better extension of existing models. This characteristic makes our object experts flexible, allowing us to replace them with different backbones.
@@ 3.2.1 Random Point Block Generation
Ground-truth block position generation. We generate the ground truth position of each random point block in the scene, which will be used to constrain the predicted position regressed by the final scene decoder. Inspired by the detection Carion et al. (2020); Misra et al. (2021); Dai et al. (2021) task, we use the 3D bounding box of each random point block as its ground truth position. By computing the mean of all points in each dimension of the entire point block, the coordinates of the center point are obtained. The half-lengths of the bounding box in each dimension are calculated by subtracting the minimum value from the maximum value in each dimension and dividing by 2. Subsequently, the center point coordinates and half-lengths in each dimension (x, y, and z) are concatenated to form the bounding box. Finally, standard procedures Misra et al. (2021) are applied to compute bounding box parameters $\bm{B_{b}}$ such as size and corners for each bounding box.
@@ 3.2.2 Object-level Block Masked Reconstruction
We first initialize our object expert model using the Point-MAE (or Point-BERT, PointGPT, etc.) model pre-trained on ShapeNet Chang et al. (2015). Since these models are open-source, we directly use the pre-trained weights provided by the official repository. This initialization allows us to incorporate object-level knowledge priors into our pipeline without the need for additional object data or specialized training.
Object point block generation. We treat each randomly selected point block in $\bm{B}$ as an object point block and transform its coordinates from the scene space to the object space for processing by the object expert. Specifically, we apply a simple set of transformation functions to each point block. First, we subtract the coordinates of the center point from the $N_{o}$ local points to obtain the relative coordinates of each point. Then, we normalize these coordinates to the range [-1, 1]. Finally, after applying a random rotation transformation to each point block, we obtain all point blocks $\bm{B_{o}=\{B_{o}^{1},...,B_{o}^{K_{o}}\}}$ as input to the object expert encoder. Through these transformation functions, we convert the coordinates of each point block from the scene space to the object space, decoupling the point block object coordinates from the original scene coordinates. This enables the object expert to learn the universal shape features of the point block objects.
Object point block masked and reconstruction. We use a shared object autoencoder to perform mask-based reconstruction self-supervised learning on all generated object point blocks $\bm{B_{o}}$, enabling our object expert to learn a general representation of objects. We illustrate the entire mask and reconstruction process for the object point blocks using Point-MAE Pang et al. (2022) as an example. For the $i$-th object point block $\bm{B_{o}^{i}\in\mathbb{R}^{N_{o}\times 3}}$, we use farthest point sampling and the K-nearest neighbors algorithm to divide it into $M_{o}$ point patches. Then, after randomly masking most of the patches, we generate initial tokens and positional encodings for each unmasked patch using MLP-based token encoding and positional encoding. By adding them, we obtain the token $\bm{E}_{o}^{0}\in\mathbb{R}^{rM_{o}\times C_{o}}$ for each unmasked patch, where $r$ represents the unmasked ratio, and $C_{o}$ denotes the object feature dimension. Finally, we use a shared object expert encoder to extract object features $\bm{E}_{o}^{n_{o}}\in\mathbb{R}^{rM_{o}\times C_{o}}$, where $n_{o}$ is the number of layers in the scene expert.
@@ 3.2.3 Scene-level Block Position Regression
Scene encoding. Given an input point cloud $\bm{P_{s}}\in\mathbb{R}^{N_{s}\times 3}$ with $N_{s}$ points, we first use farthest point sampling and the K-nearest neighbors algorithm to partition it into blocks. Then, using an MLP-based token encoding layer and a positional encoding layer, we generate the semantic token and positional encoding for each patch. By adding them, we obtain the token $\bm{E}_{s}^{0}\in\mathbb{R}^{M_{s}\times C_{s}}$ for each patch, where $M_{s}$ represents the number of scene patches, and $C_{s}$ denotes the scene feature dimension. Finally, we use a scene encoder based on the standard Transformer Vaswani et al. (2017) architecture to extract scene features $\bm{E}_{s}^{n_{s}}\in\mathbb{R}^{M_{s}\times C_{s}}$, where $n_{s}$ is the number of Transformer layers in the scene encoder.
Scene decoding and position regression. We apply max pooling to the features of all blocks output by the object encoder to obtain the global feature for each block. After passing through the projection layer, these point block features are transformed into features $\bm{B}_{g}\in\mathbb{R}^{K_{o}\times C_{s}}$ for scene decoding input. However, we stop the gradients of $\bm{B}_{g}$ from propagating backward into the mask reconstruction pipeline during the backpropagation process, thereby mitigating the multi-task interference caused by the scene regression task on the object encoder. We will provide a detailed explanation of this issue in Section 4.4.2. We then add the transformed point block features to randomly initialized queries to obtain enhanced queries $\bm{Q^{0}}\in\mathbb{R}^{q\times C_{s}}$, where $q$ is the number of queries. Since the number of point blocks and queries often differ, we replicate $\bm{B}_{g}$ to match all queries.
We use a Transformer decoder based on self-attention and cross-attention as our scene expert decoder. The input queries $\bm{Q^{0}}$, with the assistance of the encoded features $\bm{E}_{s}^{n_{s}}$, pass through our decoder to obtain the decoded query features $\bm{Q^{m_{s}}}$, where $m_{s}$ is the number of scene expert decoder. Finally, we use different MLP-based reconstruction heads to predict the 3D bounding box $\bm{B_{b}^{p}}$ of each random point block.
@@
We use a combination of mask reconstruction loss and point block regression loss to jointly constrain our pre-training process. For the reconstruction loss calculation, we follow previous work Pang et al. (2022); Zhang et al. (2022) and use Chamfer Distance Fan et al. (2017) (CD) as the loss function. For the regression loss calculation, we use Generalized Intersection over Union Rezatofighi et al. (2019) (GIoU) as the loss function. Therefore, our loss function is defined as follows:
@@
Method Reference #Params (M) ScanObjectNN ModelNet40
OBJ-BG OBJ-ONLY PB-T50-RS
Supervised Learning Only
PointNet Qi et al. (2017a) CVPR 2017 3.5 73.3 79.2 68.0 89.2
PointNet++ Qi et al. (2017b) NeurIPS 2017 1.7 82.3 84.3 77.9 90.7
SFR Zha et al. (2023a) ICASSP 2023 - - - 87.8 93.9
PointMLP Ma et al. (2022) ICLR 2022 12.6 - - 85.2 94.5
Self-Supervised Learning
Point-BERT Yu et al. (2022) CVPR 2022 22.1 87.43 88.12 83.07 93.2
Point-MAE Pang et al. (2022) ECCV 2022 22.1 90.02 88.29 85.18 93.8
Point-M2AE Zhang et al. (2022) NeurIPS 2022 15.3 91.22 88.81 86.43 94.0
PointGPT-S Chen et al. (2024) NeurIPS 2023 29.2 91.6 90.0 86.9 94.0
PointDif Zheng et al. (2024) CVPR 2024 - 93.29 91.91 87.61 -
ACT Dong et al. (2023) ICLR 2023 22.1 93.29 91.91 88.21 93.7
LCM Zha et al. (2024b) NeurIPS 2024 2.7 94.51 92.75 88.87 94.2
Point-FEMAE Zha et al. (2024a) AAAI 2024 27.4 95.18 93.29 90.22 94.5
Point-BERT (baseline) CVPR 2022 22.1 92.48 91.60 87.91 93.2
Point-MAE (baseline) ECCV 2022 22.1 92.67 92.08 88.27 93.8
PointGPT-B (baseline) NeurIPS 2023 120.5 95.8 95.2 91.9 94.4
Point-MoDE w/ Point-BERT Ours 22.1 93.46($\uparrow$ 1.0) 92.25($\uparrow$ 0.7) 88.58($\uparrow$ 0.7) 93.6($\uparrow$ 0.4)
Point-MoDE w/ Point-MAE Ours 22.1 93.98($\uparrow$ 1.3) 92.77($\uparrow$ 0.7) 89.14($\uparrow$ 0.9) 94.0($\uparrow$ 0.2)
Point-MoDE w/ PointGPT-B Ours 120.5 96.9($\uparrow$ 1.1) 95.6($\uparrow$ 0.4) 92.4($\uparrow$ 0.5) 94.7($\uparrow$ 0.3)
Table 1: Classification accuracy on real-scanned (ScanObjectNN) and synthetic (ModelNet40) point clouds. In ScanObjectNN, we report the overall accuracy (%) on three variants. In ModelNet40, we report the overall accuracy (%) for both without and with voting. ”#Params” represents the model’s parameter count. $\dagger$ indicates that the pre-trained model used simple rotational augmentation during fine-tuning on ScanObjectNN.